# HAFixAgent: History-Aware Automated Program Repair Agent

### v2

@@ 6.1.3. Results
Table 3. Effectiveness comparison of HAFixAgent against RepairAgent and BIRCH-feedback on their shared subsets. For HAFixAgent and BIRCH-feedback, cells report the number of bugs with plausible patches that pass the full test suite. For RepairAgent, we report the number of bugs with correct patches because their repository does not provide plausible bug IDs, and only reports aggregates (plausible = 186, correct = 164). Since correct is a subset of plausible, RepairAgent’s counts are more conservative compared to the plausible numbers for BIRCH-feedback.
(a) 829 common bugs with RepairAgent
Category Common HAFixAgent RepairAgent
@@
SL 165 135 134 134 138 99
SH 128 93 91 102 98 30
SFMH 408 224 238 235 239 31
MFMH 128 50 47 43 48 4
Total 829 502 510 514 523 164 (186*)
    The 186* denotes the total number of plausible patches of RepairAgent; specific bug IDs are only available for correct patches in their repository.
(b) 371 common multi-hunk bugs with BIRCH-feedback
Category Common HAFixAgent BIRCH-feedback
@@
SFMH 244 123 125 133 126 100
MFMH 127 49 46 42 47 33
Total 371 172 171 175 173 133
HAFixAgent substantially outperforms both RepairAgent and BIRCH-feedback. On the 829 bugs shared with RepairAgent (Table 3(a)), the best context configuration of HAFixAgent (fl_diff) mode repairs 523 bugs versus 164 correct fixes reported by RepairAgent, a +218.9% overall improvement. Across different context configurations of HAFixAgent, the improvements relative to this 164-fix baseline range from +206.1% to +218.9% (502, 510, 514, 523), with an average of 212.3% improvement. When comparing against RepairAgent’s reported aggregate plausible total of 186 (the equivalent metric), the corresponding improvements are +170.0% to +181.2% (175.4% by average).
On the 371 bugs shared with BIRCH-feedback (Table 3(b)), HAFixAgent also achieves higher totals. The best context configuration (fn_pair) repairs 175 bugs versus 133 for BIRCH-feedback, a +31.6% improvement. Across four context configurations, the gains range from +28.6% to +31.6%, with an average of 29.9% improvement. By category, the best HAFixAgent configuration (fn_pair) repairs 133 SFMH and 42 MFMH bugs, compared with 100 and 33 for BIRCH-feedback. Overall, HAFixAgent is consistently effective than both SOTA baselines across shared subsets and categories.
We note, however, that these SOTA comparisons are confounded by the use of different LLMs (RepairAgent used GPT-3.5, BIRCH-feedback used o4-mini, and our work uses DeepSeek-V3.2-Exp). Therefore, this analysis primarily establishes HAFixAgent’s competitiveness, while our internal ablation study (comparing history to non-history with the same LLM) provides the direct, controlled evidence for our claims.
Table 4. Repair performance across different context configurations over 4 bug categories on Defects4J. # Pass represents the number of bugs passing the test suite. Unique # Pass represents the number of bugs solved by that configuration but not by the corresponding non-history configuration. The highest value for each category group in each column is highlighted in bold.
Category Context # Pass Plausible@1 # Unique Pass
SL non-history 137 82.0% -
fn_all 136 81.4% 10
fn_pair 136 81.4% 10
fl_diff 140 83.8% 11
SH non-history 97 72.4% -
fn_all 96 71.6% 16
fn_pair 108 80.6% 19
fl_diff 103 76.9% 18
SFMH non-history 238 56.0% -
fn_all 251 59.1% 67
fn_pair 249 58.6% 69
fl_diff 254 59.8% 68
MFMH non-history 50 39.1% -
fn_all 47 36.7% 13
fn_pair 43 33.6% 13
fl_diff 48 37.5% 12
Figure 3. Venn diagram comparing bug repair performance across 4 context configurations (non-history, fn_all, fn_pair, and fl_diff)
over 4 bug categories on Defects4J.
@@
History configurations fix more unique bugs compared to the non-history configuration, with the largest gains for the multi-hunk bug categories. Table 4 summarizes the number of passing and uniquely passing bugs across the four context modes, while Figure 3 visualizes their overlap and complementarity. Across all 854 bugs, the three history-based configurations add 194 bugs that the non-history configuration never fixes, whereas the non-history configuration contributes only 32 unique fixes not recovered by any history configuration. By category, fl_diff performs best on SL with 140 total fixes and 11 unique ones and fn_pair leads on SH with 108 fixes and 80.6% Plausible@1. In SFMH, fl_diff achieves 254 fixes and the three history configurations collectively add 119 more repairs, while the non-history configuration uniquely fixes 17. Finally, the MFMH category reveals a complex trade-off. Although the non-history configuration achieves the highest total fixes (50), this aggregate score masks a clear complementarity. Collectively, the history configurations add 24 unique fixes not found by the non-history agent, outweighing the 8 unique fixes from the non-history side (Figure 3). This suggests that while history can sometimes distract the agent on complex multi-file bugs, it also provides an orthogonal signal that unlocks repairs for a different subset of bugs.
@@
The three history configurations are complementary rather than redundant, solving distinct sets of bugs that others miss. The Venn diagrams in Figure 3 show large shared cores in each category (SL 118, SH 72, SFMH 140, MFMH 22), yet each heuristic contributes nontrivial unique fixes, especially in SFMH, where fn_all, fn_pair, and fl_diff add 21, 19, and 19 unique cases, respectively. This pattern shows that fn_all captures co-evolution across functions, fn_pair captures before and after semantics, and fl_diff captures fine-grained edits at the file level, so the union of history-based configurations yields broader repair coverage than any single configuration.
Summary for RQ1: (1) HAFixAgent outperforms RepairAgent by an average of 212.3% across the different context configurations, and outperforms BIRCH-feedback by an average of 29.9%. Additionally, across all 854 bugs, history configurations add 194 history-only fixes versus 32 non-history-only, with the largest gains in multi-hunk categories. (2) The three history heuristics solve distinct cases beyond a shared core, indicating that combining heuristics yields broader coverage than any single setting.
6.2. RQ2: What Are the Cost and Efficiency of HAFixAgent?
@@ 6.2.1. Motivation
While RQ1 demonstrates that incorporating historical information into HAFixAgent enhances repair performance, it is also important to understand the associated cost and efficiency trade-offs. Cost is a first-order concern for agent systems because iterative tool use can accumulate substantial token and time budgets, even on simple bugs (bouzenia2024repairagent; xia2024automated). Recent evaluations also highlight the risk of “token snowballs” and expensive failures (a phenomenon where unresolved attempts consume several times more tokens than successful ones (DBLP:journals/corr/abs-2509-09853), making cost and efficiency as important to measure as accuracy. HAFixAgent integrates additional contextual information into each prompt, which might potentially increase inference cost and agent reasoning steps. Therefore, we analyze whether the performance gains observed in RQ2 come at a significant cost and how efficiently the agent converges toward successful repairs across bug categories.
@@ 6.2.2. Approach
We evaluate the inference cost (USD) and agent reasoning steps across all four configurations (non-history, fn_all, fn_pair, and fl_diff) over four bug categories (SL, SH, SFMH and MFMH). Inference cost is computed from input and output token usage converted to USD using official API pricing rates published for the DeepSeek-V3.2-Exp model. Efficiency is measured by the number of agent actions taken per run. A run terminates either upon a successful repair or by hitting one of the predefined limits (50 steps, \$1 USD cost, or 1-hour timeout), as defined in Section 5.4. We report cost and step metrics separately for successful and failed attempts, allowing us to isolate and analyze the average resources required for a successful repair. Finally, we analyze the success rate versus cost across configurations to characterize category-specific trade-off frontiers.
@@ 6.2.3. Results
(a) SL
(b) SH
(c) SFMH
(d) MFMH
Median values (Success / Failed)
SL SH SFMH MFMH
non-history 12 / 50 23 / 50 24 / 50 32 / 50
fn_all 14 / 50 25 / 50 27 / 50 28 / 50
fn_pair 13 / 50 24 / 50 25 / 50 31 / 50
fl_diff 16 / 50 22 / 50 29 / 50 29 / 50
Figure 4. Agent reasoning step distributions for successful and failed repairs across different HAFixAgent context configurations.
(a) SL
(b) SH
(c) SFMH
(d) MFMH
Median values (Success / Failed)
SL SH SFMH MFMH
non-history 0.005 / 0.038 0.016 / 0.047 0.016 / 0.045 0.029 / 0.038
fn_all 0.008 / 0.050 0.019 / 0.046 0.020 / 0.047 0.026 / 0.047
fn_pair 0.005 / 0.046 0.015 / 0.039 0.016 / 0.045 0.023 / 0.040
fl_diff 0.008 / 0.042 0.014 / 0.047 0.021 / 0.047 0.023 / 0.043
Figure 5. Inference cost distributions for successful and failed repairs across different HAFixAgent context configurations.
@@
Successful repairs converge well before the step cap. Figure 4 reports agent step distributions by configuration and outcome. Across categories, the median number of steps for successful repairs falls far below 50, whereas for failed attempts it sits at the cap of 50 for all configurations. The median number of steps for successful repairs is 12-16 in SL, 22-25 in SH, 24-29 in SFMH, and 28-32 in MFMH, showing that harder categories require more steps but still finish well under the cap.
@@
History does not significantly increase step counts. To rigorously assess whether historical heuristics significantly alter repair costs or the number of steps compared to the non-history configuration, we conducted Friedman tests followed by pairwise Wilcoxon signed-rank tests with Bonferroni correction ($\alpha=0.0167$) on bugs successfully fixed by all four configurations (Table 5). Within each category, the success medians for non-history, fn_all, fn_pair, and fl_diff differ by only a few steps. The Friedman tests on this matched set detect no significant step differences in any category (all $p\geq 0.05$). Notably, in MFMH, the history configurations are as fast or faster than the non-history configuration (non-history 32 vs fn_all 28, fn_pair 31, fl_diff 29), and in SH, the fl_diff configuration converges the quickest (median 22). This confirms that adding historical context does not create a significant step overhead.
Agent cost is dominated by failure, not configuration. Figure 5 summarizes inference cost distributions by configuration and outcome. Across categories, failed attempts cost markedly more than successful ones, clustering near 0.03-0.05 USD median, while successes are cheaper: SL 0.005-0.008, SH 0.014-0.019, SFMH 0.016-0.021, and MFMH 0.023-0.029. This mirrors the step analysis, where failures hit the step cap and thus consume more tokens.
Adding historical context keeps the cost comparable in most configurations. For SL, the Friedman test is significant for cost, with the Wilcoxon post-hoc tests showing that fn_all ($p=0.0022$) and fl_diff ($p=0.001$) cost more than non-history, while fn_pair shows no significant difference ($p=0.0444>0.0167$). For SH, SFMH, and MFMH, Friedman tests are not significant for cost (all $p\geq 0.05$). Notably, the median costs for successful repairs (shown in Figure 5) can be as low or lower with history in harder categories (for example, MFMH median success costs drop from 0.029 to 0.026-0.023), but these reductions are not statistically significant on the matched set (i.e., the subset of bugs fixed by all four configurations).
Figure 6. Trade-off analysis of success rate and cost per bug across different combinations of configurations.
Table 5. Statistical Comparison: History Heuristics vs. Non-history (Bonferroni $\alpha=0.0167$). The last three columns report p-values for each history configuration compared against the non-history configuration.
Category Metric N Friedman p Pairwise p-values
fn_all fn_pair fl_diff
SL Cost 118 <0.001<0.001 0.0022 0.0444 0.001
Steps 118 0.176
SH Cost 72 0.184
Steps 72 0.400
SFMH Cost 140 0.122
Steps 140 0.051
MFMH Cost 22 0.142
Steps 22 0.427
  • N = number of bugs successfully fixed by all 4 configurations.
  • Bold values indicate $p<0.0167$.
  • — indicates the Friedman test was not significant ($p\geq 0.05$); post-hoc tests were not performed.
Combining history configurations increases success but also increases cost, exhibiting diminishing returns, and the most cost-effective choices are usually two or three configurations. Figure 6 plots success rate versus average cost per bug by category, with marker shape indicating how many configurations are combined. For Single-Line bugs, fn_pair +fl_diff reaches 90.4% at an average cost of 0.041 USD, close to the union of all four configurations, 95.2% at 0.086 USD. For Single-Hunk bugs, fn_all +fn_pair attains 87.3% at 0.063, while the four-configuration union gives 94.0% at 0.130. For Single-File-Multi-Hunk bugs, fn_all +fl_diff yields 73.9% at 0.074, compared to 84.0% at 0.147 with the four-configuration union. For Multi-File-Multi-Hunk bugs, non-history +fn_pair provides 49.2% at 0.086 before the cost rises to 57.8% at 0.182 with all four. Across categories, two or three history configurations give the best success per dollar. As these results show, adding the final configuration to create the four-configuration union raises the cost substantially for only small gains, demonstrating clear diminishing returns.
Summary for RQ2: (1) Successful repairs finish well before the step cap, while failures commonly hit the cap. (2) History-based APR agents are as efficient as non-history in terms of step count and cost (except for SL). (3) Combining two or three history configurations delivers the best success, while using all configurations yields small extra gains at a significantly higher cost.
7. Discussion and Future Work
In this section, we provide a discussion of our results and introduce future work.
7.1. Performance Regressions

### v3

@@ 6.1.3. Results
Table 2. Same-LLM effectiveness comparison on Defects4J (all methods use DeepSeek-V3.2-Exp). Numbers show plausible patches passing the full test suite. Bold = best per category.
(a) All 854 Defects4J bugs: HAFixAgent vs RepairAgent
Cat. Total HAFixAgent RepairAgent
@@
SL 167 137 136 136 140 121
SH 134 97 96 108 103 68
SFMH 425 238 251 249 254 141
MFMH 128 50 47 43 48 18
Total 854 522 530 536 545 348
(b) 371 multi-hunk bugs: HAFixAgent vs BIRCH-feedback
Cat. Common HAFixAgent BIRCH-fb
@@
SFMH 244 123 125 133 126 85
MFMH 127 49 46 42 47 34
Total 371 172 171 175 173 119
HAFixAgent substantially outperforms both RepairAgent and BIRCH-feedback on the same LLM. On all 854 Defects4J bugs (Table 2(a)), the best HAFixAgent configuration (fl_diff) repairs 545 bugs versus 348 for RepairAgent, a +56.6% improvement. Even HAFixAgent-non-history (522), which uses no historical context, outperforms RepairAgent by +50.0%, demonstrating that the agent architecture itself is already more effective. The gap is largest for complex bugs: on SFMH, HAFixAgent-fl_diff repairs 254 versus 141, and on MFMH, HAFixAgent-non-history repairs 50 versus 18.
On the 371 multi-hunk bugs evaluated by BIRCH-feedback (Table 2(b)), the best HAFixAgent configuration (fn_pair) repairs 175 bugs versus 119 for BIRCH-feedback, a +47.1% improvement. By category, HAFixAgent-fn_pair repairs 133 SFMH and 42 MFMH bugs, compared with 85 and 34 for BIRCH-feedback. Since all methods now use the same LLM, these improvements directly reflect differences in repair strategy rather than model capability.
Table 3. History ablation on Defects4J (854 Java bugs) and BugsInPy (501 Python bugs). Each cell shows #Pass (Plausible@1). Superscript = #Unique bugs fixed by that config but not by non-history. Bold = best per category.
Dataset Category Total non-history fn_all fn_pair fl_diff
Defects4J SL 167 137 (82.0%) 136 (81.4%) +10 136 (81.4%) +10 140 (83.8%)+11
SH 134 97 (72.4%) 96 (71.6%) +16 108 (80.6%)+19 103 (76.9%) +18
SFMH 425 238 (56.0%) 251 (59.1%) +67 249 (58.6%) +69 254 (59.8%)+68
MFMH 128 50 (39.1%) 47 (36.7%) +13 43 (33.6%) +13 48 (37.5%) +12
Total 854 522 (61.1%) 530 (62.1%) +106 536 (62.8%) +111 545 (63.8%)+109
BugsInPy SL 82 74 (90.2%) 73 (89.0%) +4 73 (89.0%) +7 71 (86.6%) +5
SH 126 93 (73.8%) 100 (79.4%) +20 105 (83.3%)+20 93 (73.8%) +16
SFMH 202 54 (26.7%) 148 (73.3%) +107 137 (67.8%) +95 157 (77.7%)+118
MFMH 91 51 (56.0%) 56 (61.5%) +23 59 (64.8%)+17 53 (58.2%) +21
Total 501 272 (54.3%) 377 (75.2%)+154 374 (74.7%) +139 374 (74.7%) +160
Figure 3. Venn diagrams comparing bug repair overlap across 4 context configurations on Defects4J (top) and BugsInPy (bottom). Each number shows the count of bugs uniquely fixed by that intersection of configurations.
Table 4. Median cost (USD) and agent steps for successful repairs, with Friedman test p-values for cost on the matched set (N = bugs fixed by all 4 configs). Friedman tests for steps were non-significant on all categories (all $p\geq 0.05$). Bold p-values indicate significance ($p<0.05$); marks configs with significantly higher cost than non-history (pairwise Wilcoxon, Bonferroni $\alpha=0.0167$).
Median Cost (USD) Median Steps
Dataset Cat. N non-history fn_all fn_pair fl_diff non-history fn_all fn_pair fl_diff Fried. p
Defects4J SL 118 0.005 0.008 0.005 0.008 12 14 13 16 <<0.001
SH 72 0.016 0.019 0.015 0.014 23 25 24 22 0.184
SFMH 140 0.016 0.020 0.016 0.021 26 28 26 30 0.122
MFMH 22 0.029 0.026 0.023 0.023 34 29 32 30 0.142
BugsInPy SL 59 0.027 0.029 0.020 0.033 19 19 18 18 0.214
SH 61 0.030 0.031 0.034 0.036 23 18 23 21 0.200
SFMH 27 0.070 0.037 0.055 0.055 26 24 24 23 0.157
MFMH 20 0.062 0.064 0.066 0.060 29 30 31 30 0.204
@@
On Defects4J, history improves repair by +4.4% (545 vs 522), with strong complementarity across categories. Table 3 (top) shows the Defects4J ablation. Across all 854 bugs, the best history configuration (fl_diff) repairs 545 versus 522 for non-history (+4.4%). By category, fl_diff performs best on SL (140 fixes, 83.8%) and SFMH (254 fixes, 59.8%), while fn_pair leads on SH with 108 fixes (80.6%). MFMH reveals a trade-off: non-history achieves the highest total (50), yet history configurations collectively add 24 unique fixes not found by non-history, outweighing the 8 unique non-history-only fixes (Figure 3). This indicates that history provides an orthogonal signal even when it does not improve the aggregate count.
On BugsInPy, history improves repair by +38.6% (377 vs 272), with SFMH gaining +190.7% (54 to 157). Table 3 (bottom) shows the BugsInPy ablation. The best history configuration (fn_all) repairs 377 out of 501 bugs (75.2%) versus 272 for non-history (54.3%), a gain of +105 bugs (+38.6%). The SFMH category drives the largest improvement: fl_diff repairs 157 SFMH bugs (77.7%) compared with only 54 for non-history (26.7%), a +190.7% improvement. History also helps on SH (+12 bugs with fn_pair) and MFMH (+8 bugs with fn_pair). The substantially larger history benefit on BugsInPy compared to Defects4J may reflect differences in benchmark characteristics, such as language, project complexity, or the lower baseline performance of the non-history configuration on BugsInPy (54.3% vs 61.1%), leaving more room for history to contribute.
@@
The three history configurations are complementary, solving distinct sets of bugs. The Venn diagrams in Figure 3 show large shared cores in each category, yet each heuristic contributes nontrivial unique fixes. On Defects4J SFMH, fn_all, fn_pair, and fl_diff each add 21, 19, and 19 unique cases beyond the 140-bug shared core. On BugsInPy SFMH, the shared core is only 27 bugs, with history configurations expanding the fixable set dramatically (72 bugs in the three-way history overlap alone). This pattern confirms that fn_all (function-name co-evolution), fn_pair (before/after semantics), and fl_diff (fine-grained edits) capture complementary signals. As shown in the summary of Figure 3, the union of all three history configurations fixes 684 Defects4J bugs versus 522 for non-history (+31.0%), and 463 BugsInPy bugs versus 272 (+70.2%), with the largest per-category gains on SFMH and MFMH.
6.2. RQ2: Efficiency and Robustness
@@ 6.2.1. Motivation
While RQ1 shows that historical context improves repair effectiveness, two practical concerns remain. First, cost is a primary concern for agent systems, as iterative tool use can accumulate substantial token and time budgets, even on simple bugs (Bouzenia et al., 2025; Xia and Zhang, 2024). Recent evaluations also highlight the risk of expensive failures where unresolved attempts consume several times more tokens than successful ones (Fan et al., 2025). Since HAFixAgent integrates additional contextual information into each prompt, it is important to verify that the performance gains from RQ1 do not incur significant cost overhead. Second, all RQ1 experiments assume perfect fault localization (FL), which is unrealistic in practice (Yang et al., 2024; Fan et al., 2023). In practice, FL tools produce imprecise locations, and when blame is computed on incorrect lines, irrelevant commits may be retrieved, potentially negating the history benefit. This RQ addresses two concerns: (a) whether adding historical context inflates cost, and (b) whether the history benefit persists under noisy FL.
@@ 6.2.2. Approach
We evaluate inference cost (USD) and agent reasoning steps across all four configurations on Defects4J (854 bugs) and BugsInPy (501 bugs). Inference cost is computed from token usage converted to USD using API pricing rates for DeepSeek-V3.2-Exp. A run terminates upon a successful repair or by hitting one of the predefined limits (50 steps, \$1 USD cost, or 1-hour timeout), as defined in Section 5.4. To ensure fair comparison, we report cost and step metrics on the subset of bugs fixed by all four configurations, and assess statistical significance using Friedman tests followed by pairwise Wilcoxon signed-rank tests with Bonferroni correction.
To evaluate robustness under imperfect FL, we simulate noise by applying positive line shifts of +1, +3, and +5 to the fault locations of a stratified sample of 100 Defects4J bugs (25 per category: SL/SH/SFMH/MFMH, randomly sampled with seed=42, excluding the Chart project). The shift affects the entire pipeline: the agent receives shifted locations in its prompt, blame is computed on the shifted lines (potentially retrieving different or irrelevant commits), and history context is extracted from the shifted blame. The existing perfect-FL results from RQ1 serve as the shift=0 baseline. We run all 4 configurations at each shift level, yielding $100\times 4\times 3=1{,}200$ additional runs. Running the full datasets across all shifts would require 16,260 runs and is prohibitively expensive, so we adopt stratified sampling with equal representation across bug categories to maintain coverage while keeping costs manageable.
@@ 6.2.3. Results
Figure 4. Success rate degradation under noisy fault localization (line shifts +1, +3, +5) on 100 Defects4J bugs (25 per category). Shift=0 uses perfect FL from RQ1.
@@
Successful repairs converge well before the step cap. Table 4 reports median agent steps and cost by configuration and category on both datasets. On Defects4J, the median steps for successful repairs is 12 to 16 in SL, 22 to 25 in SH, 26 to 30 in SFMH, and 29 to 34 in MFMH, all well below the 50-step cap. BugsInPy shows a similar pattern (18 to 19 in SL, 18 to 23 in SH, 23 to 26 in SFMH, 29 to 31 in MFMH). Harder categories require more steps but still finish well under the cap on both benchmarks.
@@
History does not significantly increase step counts or cost. To rigorously assess whether historical heuristics significantly alter repair costs compared to non-history, we conducted Friedman tests on the matched set of bugs fixed by all four configurations (Table 4, Friedman p column). On Defects4J, only SL shows a significant cost difference ($p<0.001$); pairwise Wilcoxon tests with Bonferroni correction ($\alpha=0.0167$) show that fn_all ($p=0.0022$) and fl_diff ($p=0.001$) cost slightly more than non-history on SL. For SH, SFMH, and MFMH, costs are statistically comparable (all $p\geq 0.05$). Notably, the median costs for successful MFMH repairs are comparable or lower with history (e.g., Defects4J: non-history \$0.029 vs fn_pair \$0.023; BugsInPy: non-history \$0.062 vs fl_diff \$0.060).
The same cost pattern holds on BugsInPy. Table 4 reports median cost and steps for both datasets, along with Friedman test p-values. On Defects4J, only SL shows a significant cost difference ($p<0.001$); for SH, SFMH, and MFMH, costs are statistically comparable across configurations. On BugsInPy, no category shows significant differences (all $p\geq 0.15$), confirming that history does not inflate cost on either benchmark. Notably, on BugsInPy SFMH, the non-history configuration is the most expensive (\$0.070 median) while history configurations cost \$0.037 to \$0.055, suggesting that history helps the agent converge faster on complex bugs.
Combining history configurations increases success but also increases cost, with diminishing returns. On Defects4J, two-config combinations such as fn_pair +fl_diff reach 90.4% on SL at \$0.041, close to the four-config union (95.2% at \$0.086). On BugsInPy, the pattern is similar but with higher absolute costs. Across both datasets, two or three configurations give the best success per dollar, while adding the fourth yields small extra gains at substantially higher cost.
Under noisy FL, history provides increasing resilience: on SFMH, non-history collapses to 0% at shift=+5 while history maintains 40 to 56%. Figure 4 shows success rates as FL noise increases. The most pronounced effect appears on SFMH: the non-history configuration drops from 64% at shift=0 to 0% at shift=+5. In contrast, history configurations maintain 40 to 56% success at shift=+5 (fl_diff and fn_all both at 56%), indicating that historical context acts as a safety net when FL is imprecise. The history advantage on SFMH grows with noise, from 4-4% at shift=0 to +56+56% at shift=+5.
On SH, a similar pattern appears at moderate noise: at shift=+3, the best history configurations (fn_all and fn_pair, both 88%) outperform non-history (68%) by 20%. At shift=+5, all configurations degrade, but history still leads (36% vs 24%). SL bugs remain robust across configurations, consistent with their simplicity, while MFMH shows no clear trend due to the small sample size and inherent difficulty. Although individual data points may fluctuate with 25 bugs per category, the overall degradation trend and the widening gap between history and non-history remain consistent. These results show that history-aware repair not only benefits from precise fault locations but also provides contextual value that becomes increasingly important as FL precision decreases.
7. Discussion